### v3

# Prompting Large Language Models with Divide-and-Conquer Program for Discerning Problem Solving

Abstract
Foundation models, such as Large language Models (LLMs), have attracted significant amount of interest due to their large number of applications. Existing works show that appropriate prompt design, such as Chain-of-Thoughts, can unlock LLM’s powerful capacity in diverse areas. However, when handling tasks involving repetitive sub-tasks and/or deceptive contents, such as arithmetic calculation and article-level fake news detection, existing prompting strategies either suffers from insufficient expressive power or intermediate errors triggered by hallucination. To make LLM more discerning to such intermediate errors, we propose to guide LLM with a Divide-and-Conquer program that simultaneously ensures superior expressive power and disentangles task decomposition, sub-task resolution, and resolution assembly process. Theoretic analysis reveals that our strategy can guide LLM to extend the expressive power of fixed-depth Transformer. Experiments indicate that our proposed method can achieve better performance than typical prompting strategies in tasks bothered by intermediate errors and deceptive contents, such as large integer multiplication, hallucination detection and misinformation detection.
Machine Learning, ICML
1 Introduction
Large language models (LLM) based on the Transformer architecture have led to major breakthroughs in natural language processing and other related fields in artificial intelligence(Brown et al., 2020; Radford et al., ; Touvron et al., 2023). State-of-the-art general-purpose language models have demonstrated remarkable advancements in various domains, including question answering, graph learning, reading comprehension, text generation, and machine translation (Chen et al., 2023b; Tan et al., 2023; Hendy et al., 2023; Mao et al., 2023; Zong & Krishnamachari, 2023). These developments paves the way towards general-purpose problem solvers (Bubeck et al., 2023).
However, as pointed out in (Wei et al., 2022), significant challenges arise when scale-up models are applied to tasks involved with long solution paths, such as those requiring mathematical or knowledge reasoning. A series theoretic works attribute this challenge to Parallelism Tradeoff (Merrill & Sabharwal, 2023a), a fundamental limitation of Transformers. Specifically, unlike Recurrent Neural Network whose computational depth is linear to the input sequence length (i.e., the depth is $O(n)$, where $n$ is the input sequence length), Transformer does not contain any recurrent structure. Such design, while achieving superior parallelizability than RNN, makes Transformers suffer from limited expressive power. Merrill & Sabharwal proved that the expressive power of fixed-depth log-precision Transformer, which is very close to the most commonly applied Transformer architecture for LLMs, is bounded by constant-depth logspace-uniform threshold circuits. Thus, they fail to accurately tackle the tasks requiring long solution paths.
Figure 1: An illustrative example of hallucination detection with entangled problem solving (i.e., directly forward all inputs into the LLM) and disentangled problem solving (i.e., divide the problem inputs to parallel sub-tasks and tackle them parallelly). The sentence marked with red back font in the material is the evidence that contradict with the first claim in summary (marked with red font).
To address this challenge, carefully designed prompting strategies have been developed to tackle tasks that requires stronger expressive power (Feng et al., 2023). A series of works focus on prompting the LLM to output the intermediate steps that derive the final answer in an autoregressive manner such as Chain-of-Thoughts (CoT) (Wei et al., 2022; Wang et al., 2022; Yao et al., 2023; Zhou et al., 2022; Chen et al., 2023a). Theoretically, these prompting strategies convert the role of Transformer from the complete problem solver to a sub-problem solver in a dynamic programming or tree searching algorithm (Merrill & Sabharwal, 2023b). In this way, these prompting strategies expand the expressive power of the LLMs and successfully improve the reasoning and searching ability of LLMs (Feng et al., 2023). However, in most of these aforementioned prompting strategies, the processes of sub-problem decomposition, sub-problem resolution, and sub-resolution assembly are intertwined together during the autoregressive token decoding. Such design makes the sub-problem generation process lack control from the human or agent side and susceptible to disruptions of the task resolution process. As a result, when the task requires a large quantity of repetitive tasks (e.g. large integer multiplication and article-level verification), LLM is prone to intermediate errors, such as missing some inputs or generating wrong sub-tasks, leading to problematic answers (Amaro et al., 2023). This phenomenon is especially serious when the task input contains deceptive information or contents that could trigger hallucination (Chen & Shu, 2023; Li et al., 2023a). An illustrative example is presented in the Fig. 1. To alleviate this issue, some program guided prompting strategies such as Least-to-Most (Zhou et al., 2022) (LtM) and Decomposed Prompting (Khot et al., 2022) propose to disentangle sub-task generation and resolution. However, they implement the task decomposer through multi-round conversation (or question-answering) with a LLM to sequentially raise and solve the sub-problems in an alternate manner. When tackling deceptive text like hallucination and fake news, such intertwined process often guides the LLM to sequentially tackle the corpus and thus follow the context’s flow, making LLM prone to deception.
In fact, human brains also suffer from similar hallucination issues (Liu & Sajda, 2023), especially when the tasks are too hard or too complex. For example, when reviewing a long academic paper, some reviewers produce low-quality reviews (Garcia et al., 2021; Tennant & Ross-Hellauer, 2020; Cortes & Lawrence, 2021) containing hallucination-like intermediate errors, such as pointing out some ‘missing baselines’ that have already been sufficiently discussed by authors and requiring the authors to conduct ablation studies for non-existent sub-modules. To avoid such ridiculous mistakes, experienced reviewers usually think slowly (Kahneman, 2011) to follow a Divide-and-Conquer paradigm to handle this task. Specifically, they decompose the paper review as examinations of multiple central opinions. Then they retrieve supporting corpus to verify them respectively. Similar strategies are also applied in automated evaluation to the quality of text generated by LLMs (Cui et al., 2023).
Inspired by the aforementioned human experience, in this paper, we explore to guide LLM with Divide-and-Conquer program to unlock the LLM’s ability to handle tasks with repetitive sub-tasks. This strategy breaks down the whole task resolution process into three distinct sub-processes: task decomposition, sub-task resolution, and solution merge. The task decomposition process will prompt the model to separate the whole task to multiple parallelly solvable sub-tasks (i.e., solving one sub-tasks does not require the result of other sub-tasks) and list them explicitly recursively. After that the sub-task resolution process prompts the LLM to output the answer for all sub-tasks. Finally, the solution-merge process assembly the solutions recursively along the decomposition path. The above sub-processes follow a key principle that every upstream sub-process (e.g. sub-task resolution) only forward the final answer to its downstream sub-processes (e.g. solution merge) and does not forward its input and intermediate steps. We denote this principle as disentangled-sub-process principle. An illustrative figure that explain the difference between our method and previous works is provided in Fig. 2
Figure 2: The comparison between our method and the existing methods for prompting. The ellipse marks represent sub-tasks, the right-angled rectangles represent sub-task solutions, and the rounded rectangles represent intermediate steps that entangle sub-task and sub-solutions. The different shades in Tree of Thoughts (subfigure D) indicate the rates of different search directions. In CoT (Chain-of-Thoughts), CoT-SC and ToT, the Large Language Models must simultaneously generating and resolving sub-tasks. Least-to-Most (also Decomposed Prompting) disentangle sub-task generation and resolution. However, as shown in the figure, its sub-task resolution process and resolution assembly process are intertwined as it sequentially attach new sub-tasks onto the previous resolution. Different from all of them, our method totally disentangle the sub-task generation, sub-task resolution, and resolution assembly process.
To validate the expressive power of our proposed prompting strategy, we provide a theoretic analysis to validate that the proposed strategy can expand the expressive power of fixed-depth log-precision Transformer. To further empirically validate the advantage of our proposed method, we evaluate our method and representative baselines on three tasks that are challenging to existing prompting strategies even on state-of-the-art LLMs: Large Integer Multiplication, Hallucination Detection, Article-level Fact Verification (Cheng & Zhang, 2023; Li et al., 2023a; Wadden et al., 2020; Hu et al., 2023; Wu et al., 2023). These tasks either require very long reasoning paths (e.g. large integer multiplication) or contain deceptive contents (e.g. hallucination detection and fact verification), making existing methods like Chain-of-Thought prompting prone to missing intermediate steps. Meanwhile, these tasks share a common nature that the correctness of the final answer substantially relies on the thoroughness of all steps. Our experimental results show that the proposed method outperforms the baselines on all three tasks.
2 Related Work
2.1 Limitations of Transformer in Expressive Power
As discussed in previous works (Merrill & Sabharwal, 2023a; Feng et al., 2023), the expressive power of fixed-length log-precision transformers, which are widely applied in modern Pre-trained Large Language Models, is actually much more limited than people’s expects. Merrill & Sabharwal give a theoretic proof that the expressive power of fixed-length log-precision transformers is upper-bounded with ${\sf TC^{0}}$. Feng et al. further extend their analysis to explain that a lot of common problems, such as calculating arithmetic expressions, exceeds the expressive power of fixed-length log-precision transformers. Such results explains why the powerful Large Larnguage Models may make some ridiculous mistakes and why CoT can improve the performance.
2.2 Prompting Strategies for Large Language Models
In this sub-section, we introduce the existing prompting and discuss their limitations and drawbacks. Following the notations in (Yao et al., 2023), we denote the Large Language Models with parameter $\theta$ as $p_{\theta}$ and use lower case letters $x,y,z$ to denote input sequence, result, and intermediate steps, respectively.
Input-Output (IO) Prompting is the standard prompting strategy that attach input $x$ with instructions and/or few-shot in-context-learning examples to aqcuaire a prompt, denoted as ${\sf prompt}(x)$ (Yao et al., 2023). The LLM takes ${\sf prompt}(x)$ as input and predict result, i.e. $y\sim p_{\theta}(y|{\sf prompt}(x))$. The main drawback of IO prompting is its poor expressive power. Existing works have shown the expressive power limitation of IO prompting. (Merrill & Sabharwal, 2023a) proves that log-precision transformer based LLM equipped with prompt of polynomial length is a subset of ${\sf TC^{0}}$ class.
Chain-of-Thought (CoT) Prompting (Wei et al., 2022) aims at simulating human’s thinking process that handles complicated task (e.g. combinational reasoning and mathematical calculation) in a step-by-step manner. More specifically, the LLM is guided to output a series of intermediate steps $z_{1},z_{2},...,z_{n}$ (also known as thoughts) autoregressively, i.e. $z_{i}\sim p_{\theta}(z_{i}|{\sf prompt}(x),z_{1},...,z_{i-1})$. Then the LLM output the prediction of result $y$ based on the thoughts, i.e. $y\sim p_{\theta}(z_{i}|{\sf prompt}(x),z_{1},...,z_{n})$.
Exploration-of-Thought (EoT) Prompting includes a series of CoT’s variants, such as Self-consistency with CoT (CoT-SC) prompting (Wang et al., 2022) and Tree-of-Thoughts (ToT) prompting (Yao et al., 2023), which aim at addressing the limitation of CoT in exploration. Their common central idea is to generate multiple chains of thought through sampling or proposing prompting and then ensemble them to acquire a final prediction.
Program-guided Prompting aims at controlling the LLM’s generation process with symbolic programs or pre-defined procedure (Zhu et al., 2022; Jung et al., 2022; Zhou et al., 2022; Khot et al., 2022; Creswell & Shanahan, 2022; Gao et al., 2023). Among them, the Least-to-Most (LtM) Prompting (Zhou et al., 2022) and Decomposed Prompting (Khot et al., 2022) are close to this work. They are the earliest attempts that explicitly prompt the LLM to decompose the task as a series of sub-tasks and sequentially tackle them. LtM prompt a LLM to iteratively raise sub-tasks and sequentially solve them to acquire the final resolution. Decomposed Prompting can regarded as a upgraded version of LtM. It introduces special notations into the prompt to represent program states and thus can call itself (i.e., recursion) or other modules (i.e., hierarchical decomposition), endowing it stronger expressive power. Such design increased the compositional generalization ability of LLMs in different areas, such as symbolic manipulation and multi-hop QA (Khot et al., 2022).
The aforementioned CoT and EoT families incorporate LLM with stronger expressive power than IO prompting. However, a critical issue of them is that, they could miss or ignore some important intermediate steps or contents (Liu et al., 2023). This problem is even worse when we are handling tasks involved with long input (e.g. long documents and large numbers). Typical examples include large number arithmetic calculation and fact verification in long documents. Compared to them, Least-to-Most prompting and Decomposed Prompting introduce explicit task decomposition to enumerate sub-tasks. However, their task decomposers are based on multi-round conversation or question-answering, which often guides the LLM to sequentially tackle the sub-tasks. When tackling contents containing deceptive contents, e.g. hallucination and fake news detection, such design has two issues. First, it will navigate the LLM through the deceptive content’s flow and thus make LLM prone to deception. Second, as pointed by previous works (Yang & Ettinger, 2023; Dziri et al., 2024), when the round of conversation increases, the hallucination/deception will propagate and accumulate in the contexts, leading to decreased performance and accuracy.
3 Proposed Method
To avoid the task decomposition and task resolution from interweaving and interrupting each other, we propose to guide LLM with Divide-and-Conquer (DaC) program that consists of three distinct stages: task decomposition stage, sub-task resolution stage, solution merge stage. In task decomposition stage, the LLM is prompted to explicitly decompose the task as a series of parallel homogeneous sub-tasks with smaller problem sizes (e.g. divide a long paragraph to sentences). Such parallel design ensures LLMs can tackle sub-tasks without relying on a specific resolution order. As a result, the correctness of each sub-task’s resolution does not rely on other sub-tasks, making LLMs less prone to intermediate errors and deceptions, and thus the issues of multi-round conversation in LtM and Decomposed Prompting can be avoided. After that, in sub-task resolution stage, the LLM is prompted to provide the solutions for every sub-task. Finally, in the solution merge stage, the LLM is prompted to assembly the solutions of subtasks and acquire the final answer. In this process, all three stages are isolated to avoid interruption. They are all guided by a program rather than a LLM to avoid hallucination or the deception from input context. To tackle tasks of different sizes, we propose two variants: Single-Level DaC Solver and Multi-Level DaC Solver.
Algorithm 1 Single-Level Divide-and-Conquer Solver $T(S,a,t,L,f)$
0:  Input Sequence $S$, Prompt $m$ (for solution merge), Prompt $t$ (for sub-task tackling), Prompt $d$ (for task decomposition), LLM $L$
0:  Results of the task on input sequence $S$
1:  $\{S_{1},S_{2},...,S_{k}\}\leftarrow L(d,S)$
[…]
3:  for $i=1,2,...,k$ do
4:     Result $\leftarrow$ Result $+[SEP]+L(t,S_{i})$
5:  end for
6:  Return $L(m,Result)$
Single-level Divide-and-Conquer Solver decomposes the task in one call to the LLM, which expands the original task as a tree of one level. The algorithm is presented in the Alg. 1. The advantage of this variant is its simplicity and efficiency. However, when the original input is too long, single-level Divide-and-Conquer Solver may acquire sub-tasks with large problem sizes that will still trigger intermediate errors. In such a case, following (Khot et al., 2022; Lee & Kim, 2023), we can recursively expand the task as a multi-level tree. More specifically, we repeat the aforementioned steps to further divide the sub-tasks hierarchically until they are easy enough to be handled by the LLM. This can be done through a recursion program as presented the Alg. 2.
Algorithm 2 Multi-Level Divide-and-Conquer Solver Recursion $T(S,m,t,d,f,n,L)$
0:  Input Sequence $S$, Problem Size Metric Function $f(\cdot)$ (a function that measure the problem size), hyper-parameter $w$, Prompt $m$ (for merge), Prompt $t$ (for sub-task tackling), Prompt $d$ (for task decomposition), Large Language Model $L$
0:  Results of the task on input sequence $S$
1:  $S_{1},S_{2},...,S_{k}\leftarrow L(d,S)$
[…]
3:  for $i=1,2,...,k$ do
4:     if $f(S_{i})>w$ then
5:        Result $\leftarrow$ Result $+[SEP]+T(S_{i},m,t,d,f,w,L)$
6:     else
7:        Result $\leftarrow$ Result $+[SEP]+L(t,S_{i})$
8:     end if
9:  end for
10:  Return $L(m,Result)$
4 Theoretic Analysis on the Proposed Method
In this section, we provide a theoretic analysis to indicate that the proposed Divide-and-Conquer program-guided problem solving strategy help expand the expressive capacity of Transformers. Specifically, we exploit 2-color Binary Subtree Isomorphism (2-BSI) problem for the proof. 2-BSI problem is a ${\sf NC^{1}}$-complete problem. Under a widely-accepted assumption in parallel computing, we theoretically prove two facts: (1) Log-Precision Transformers with standard IO-prompting strategy can not solve 2-BSI problem, and (2) there exists a Log-Precision Transformers with Divide-and-Conquer program-guided problem solving strategy that can solve 2-BSI problem. In this way, we prove that Divide-and-Conquer strategy expand the expressive power of Large Language Models.
4.1 Fixed-Depth Transformer can not solve 2-color Binary Subtree Isomorphism (2-BSI)
In this section, we will present a problem that exceeds the expressive power of fixed-length log-precision transformers: 2-color Binary Subtree Isomorphism (also known as tree matching), an important algorithmic problem which many tasks such as semantic similarity analysis, text matching and structured text database querying can be reduced to (Kilpeläinen & Mannila, 1992; Marsi & Krahmer, 2010; He et al., 2018). More specifically, we first gives a theoretical proof that 2-color Binary Subtree Isomorphism is not solvable for fixed-length log-precision transformers.
Definition 1.
2-color Binary Subtree Isomorphism problem is that, given a pattern 2-color1112-color means that each node in the tree can be of two colors. We can understand it as that each node has a 0-1 label binary tree $t_{p}$ and a base 2-color binary tree $t_{b}$, a solver is required to judge whether the pattern tree is isomorphic to a sub-tree of $t_{b}$
In (Jenner et al., 2003), the authors pointed out that the encoding of the problem will influence the hardness of the problem. In this paper, we focus on pointer list encoding of 2-BSI. Detailed information about the pointer list encoding of 2-BSI can be found in Appendix. For pointer list encoding of 2-BSI, we have the following theorem:
Theorem 4.1.
Assume that $TC^{0}\not=NC^{1}$. For any depth $L$, any polynomial $Q$, there exists a size $n$ of pattern tree such that there exist no log-precision Transformer with a depth of $L$, hidden dimension $d<Q(n)$, and fixed prompt $p$ that can directly output the solution (Yes or No) of the 2-color Binary Subtree Isomorphism problem (2-BSI).
Proof.
The core point of the proof is that, based on (Merrill & Sabharwal, 2023a), we know that the expressive power of log-precision Transformer is in $TC^{0}$. As proved by (Jenner et al., 2003), any $L$ problem ($NC^{1}\subseteq L$) can be converted to a 2-color binary tree isomorphism problem (2-BTI, comparing whether two trees are exactly isomorphic or not) under $AC^{0}$ reduction. Therefore, 2-color binary tree isomorphism is $NC^{1}$ hard. Because any solver of 2-BSI can solve 2-color binary tree isomorphism by simply adding a layer to check whether the two trees have same size, 2-BSI is harder than 2-BSI. Since $TC^{0}\subseteq NC^{1}$ and we assume that $TC^{0}\not=NC^{1}$, then for any fixed-length polynomial-parameter log-precision transformer, there exist a size of the 2-BSI problem such that the transformer can not handle. ∎
4.2 Fixed-Depth Transformer guided by Divide-and-Conquer program solves 2-BSI
Theorem 4.2.
There exists a log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition).
Proof Sketch: The detailed proof to this theorem is provided in the Appendix A.1. Here we give a brief flow of the proof. To prove this theorem, we first show that there exists an problem size function $f(\cdot)$, hyper-parameter $w$, merge function $m(\cdot)$, sub-task tackling function $t(\cdot)$ and decomposition function $d(\cdot)$ that can solve the problem with divide-and-conquer strategy. Then we will prove that there exists one log-precision transformer with fixed depth $L$ and hidden dimension $d$ that can express $a(\cdot)$, $t(\cdot)$ and $d(\cdot)$ respectively with different but fixed-length prompts. In this way, we can prove the theorem.
5 Experiments
We evaluate the capacity of our prompting strategies on three kinds of tasks that are troubled by missing intermediate steps due to long inputs, even for state-of-the-art Language Models (e.g. GPT-3.5 and GPT-4): Multiplication and Addition of Large Numbers, Hallucination Detection in Long Context, and Fact-Checking for Misinformation Detection (Yang et al., 2023; Li et al., 2023a; Wadden et al., 2020). With these experiments, we present how Divide-and-Conquer (DaC) prompting handles this task better and suggest DaC could be a promising new paradigm in handling long documents or large inputs.
5.1 Multiplication of Long Integers
General-purpose Large Language Models, e.g. ChatGPT-3.5 and ChatGPT-4, have been bothered by its poor performance on arithmetic of big numbers, even though their expressive capacity has been proved to be enough for accurately calculate large number multiplication (Yang et al., 2023). In this experiment, we show that under zero-shot long integer multiplication, DaC can outperform other baseline prompting strategies.
Figure 3: Accuracy of our proposed method and different baseline prompting strategies on GPT-3.5 and GPT-4.
Task Setup: For this task, we randomly generated 200 pairs of 5-digit integers. We choose 5 for the digit length because according to previous works, ChatGPT-3.5 gets 0% accuracy on 4-digit multiplications (Cheng & Zhang, 2023), and ChatGPT-4 gets close to 0% accuracy on 5-digit multiplications (Yang et al., 2023). We evaluate the performance of calculation with two metrics: Accuracy and Edit Distance (Marzal & Vidal, 1993; Schaeffer et al., 2023). For accuracy, we count the result as correct one only every digit of it matches with the ground-truth answer. For edit distance, we calculate the minimum number of operations required to transform the output to the ground-truth.
Figure 4: Edit distances of DaC and other strategies on GPT-3.5 and GPT-4. Lower edit distance indicate better performance.
Strategies GPT-3.5-Turbo GPT-4
F1 Acc Prec Recall F1 Acc Prec Recall
IO-prompting 61.69 61.27 62.11 61.28 64.07 72.66 93.41 48.76
Chain-of-Thoughts 46.85 64.26 91.36 31.50 71.05 76.10 90.08 58.66
CoT-SC 47.70 64.25 88.83 32.60 71.39 76.36 90.41 58.98
Tree-of-Thoughts 70.40 59.91 55.83 95.34 69.41 71.73 75.53 64.28
Least-to-Most 56.43 64.91 74.42 45.44 72.51 77.11 90.74 60.38
Divide-and-Conquer w/o DP 67.66 67.51 67.32 68.03 73.95 76.63 83.54 66.33
Divide-and-Conquer 74.84 75.55 77.41 72.03 76.92 78.99 85.36 70.01
Table 1: Performance of different prompting methods on HaluEval dataset. We report the F1 score, Accuracy, Precision and Recall.
Setup of baselines and our method: In this task, our baselines include IO prompting, Chain of Thought (CoT), CoT-SC, Least-to-Most (LtM), and Decomposed Prompting (DeP). We also include our method’s variant: DaC without Disentangled-sub-process Principle (DaC w/o DP) for ablation study. Tree-of-Thoughts is not applicable. This is because that multiplication is deterministic calculation without requiring search in a tree. For our method, we apply Multi-Level Divide-and-Conquer program-guided solver. We set $f(a,b)$ as the minimum length of the two integers $a,b$. We set $w=2$ so that the recursion returns when both integers are no longer than $2$. If both integers are shorter, then the algorithm return. The prompts for decomposition asks the model to separate the two $n$-digit numbers to four $\frac{n}{2}$-digit numbers (e.g., separate $a,b$ as $a_{1},a_{2},b_{1},b_{2}$). The sub-task tackling prompt asks the model to calculate the multiplication of the four pairs of $\frac{n}{2}$-digit numbers and solution merge prompt merge them with bit shift and addition.
Results: Experimental results are shown in Fig. 3 and 4. As we can see, under all settings, our proposed prompting strategy outperform all the baselines. This indicate the advantage of our proposed method. Specifically, the thought-based methods, i.e. IO-Prompting, Chain-of-Thoughts and CoT-SC, all get accuracy close to 0 for both GPT-4 and GPT-3.5. The Least-to-Most prompting acquires better results on accuracy for GPT-3.5 and GPT-4 due to its advantage in disentangling task decomposition and task resolution. However, compared to our methods that can achieve accuracy better than 40% for both GPT-3.5 and GPT-4, these baselines significantly perform bad. For edit distance metric, the results are similar. For both GPT-3.5 and GPT-4, our method’s distances are lower than half of the baselines. Although for GPT-4, the Least-to-Most and DeP get edit distances that are significantly better than other baselines, it is still not comparable to our method. We claim that the advantage is brought by our method’s property in disentangling the task decomposition, task resolution and resolution assembly. The thought-based methods like CoT and CoT-SC intertwined the task decomposition and task resolution. As a result, its sub-task generation process is prone to be interrupted and gets wrong sub-task. Least-of-Most disentangle the the task decomposition and task resolution, and thus acquires much better results. However, its task resolution and resolution assembly processes are still intertwined and make the model prone to mistakes when assemblying the sub-task results. Also, DaC outperforms DaC w/o DP, indicating the importance of disentangled-sub-process principle.
5.2 Hallucination Detection in Long Context
Although Large Language Models have achieved impressive performance on various NLP tasks, they are bothered by hallucination problem (Manakul et al., 2023), especially when the generated content or the input context is too long for the user to have a thoroughly review (Zhang et al., 2023). In this paper, we focus on evaluating the performance of different strategies in guiding LLM to recognize inconsistency between given context and model response with hallucination.
Task Setup: We use the HaluEval-Summary dataset. It is one of the three datasets in HaluEval benchmark for hallucination detection, which contains the hallucination generated by ChatGPT-3.5. HaluEval-Summary have the longest context and generated contents among all three tasks in this benchmark (Li et al., 2023a). Thus, detecting hallucination on this dataset requires repeatedly verify each sentence in the response, making standard prompting strategies acquire the worst accuracy across all three tasks. We report the Accuracy, F1 score [^2] We consider the hallucination pairs as positive samples., Precision score and Recall score.
Setup of baselines, ablation variants and our method: In this task, our baselines include IO prompting, Chain of Thought, CoT-SC, Tree-of-Thoughts Least-to-Most, and Decomposed Prompting. We also include our method’s variant: DaC without Disentangled-sub-process Principle (DaC w/o DP) for ablation study. In this task, the sub-tasks are verifying fragments of the summary, which are homogeneous and do not require recurssion. In such a setting, Decomposed Prompting is equivalent to LtM. For this task, we apply single level Divide-and-Conquer solver to decompose the summary to multiple sentences, handle them separately and then merge the conclusions of all sentences. The detailed prompts are provided in the Appendix.
Results: Experimental results are shown in Tab. 1. As we can see for all other baselines, the performance on GPT-3.5 is substantially worse than GPT-4, as the hallucinations are generated by GPT-3.5 and thus more deceptive to GPT-3.5. However, for our method, when we replace GPT-4 with GPT-3.5, although the performance still decrease, the drop is significantly lower than all other baselines, indicating the robustness of our proposed method toward deceptive contents. Also, for both GPT-3.5 and GPT-4, our proposed prompting strategy outperform the baselines, presenting the advantage of our method. More specifically, compared to IO-prompting, our model achieved better performance in general, indicating the advantage brought by stronger expressive power. Meanwhile, compared to CoT and CoT-SC results, our method clearly achieved much better recall. Tree-of-Thoughts, benefited by its searching ability, acquired significantly better recall score compared to other baselines. However, its significantly lower precision substantially harm its overall performance and leads to accuracy that is even worse than standard IO-prompting. Least-to-Most, which explicitly decomposing the task and enumerating all sub-tasks, achieves better recall compared to CoT and CoT-SC. However, it follows the flow of the hallucination, leading to proneness worse recall compared to Tree-of-Thoughts and our method. In contrary, our method carefully checked all sentences, locate the one containing factual error and merge the answers. Thus, our method balance the overall performance (F1 and Accuracy score). In addition, as we can see, although DaC w/o DP outperforms other baselines in general, it still perform worse than DaC. This phenomenon indicates the importance of disentangled-sub-process principles. Meanwhile, it is noticeable that the DaC w/o DP can be regarded as the upgrade version of LtM/DeP that replace sequential sub-tasks with parallel sub-tasks. Thus, the experimental results suggest that parallel sub-task strategy makes LLM more discerning for hallucination detection.
5.3 Fact-Verification for Misinformation Detection
The increasing abuse of misinformation toward manipulating public opinions on social media has been observed in different areas, such as healthcare (e.g. the recent COVID-19 pandemic) (Sharma et al., 2020, 2022). This threat is increasingly serious due to LLM’s capacity in content generation (Li et al., 2023b; Weidinger et al., 2021; Zhang et al., 2022). This challenge raise the importance of fact-verification, which aims at judging the authenticity of an article based on a collection of evidence from verified source (Whitehouse et al., 2022; Zhang & Gao, 2023). In this experiment, we present that DaC can outperform other baselines in fact-verification involved with news article.
Strategies GPT-3.5-Turbo GPT-4
F1 G-Mean Prec Recall F1 G-Mean Prec Recall
Io-Prompting 72.12 72.77 83.22 63.64 69.15 71.77 94.44 54.55
Chain-of-Thoughts 56.09 60.64 90.48 40.64 74.03 75.79 94.21 60.96
CoT-SC 56.83 61.44 91.67 41.18 70.09 73.45 100.0 53.95
Tree-of-Thoughts 69.91 73.30 53.74 100.0 77.34 78.00 88.89 68.45
Least-to-Most 54.08 54.15 51.46 56.99 73.56 74.25 85.21 64.71
Divide-and-Conquer w/o DP 61.32 63.66 83.81 48.35 74.15 75.68 92.68 61.79
Divide-and-Conquer 76.88 77.13 83.65 71.12 81.11 81.24 76.67 86.10
Table 2: Performance of different prompting methods on SciFact dataset. We report the F1 score, G-Mean score, Precision and Recall.
Task Setup: In this experiment, we mainly adopt SciFact dataset (Wadden et al., 2020). In SciFact dataset, each sample is a pair of news and evidence, where the evidence is the abstract of a peer-reviewed paper and the news is a claim summarized by human annotators from fake news or true news. To better simulate the real-world scenario where news usually appears as an article, following Chen & Shu, we generate a dataset of article-level misinformation based on SciFact dataset. Specifically, for a given claim, we apply Large Language Model (i.e., ChatGPT-4) to extend the claim as an article based on the evidence. For this task, similar as hallucination detection, we apply single level Divide-and-Conquer solver to decompose the news article to multiple sentences, handle them separately and then merge the conclusions of all sentences. Also, the baselines in this experiments are the same as Hallucination Detection, for the same reason, Decomposed Prompting is equivalent to LtM. The evaluation metrics includes F1 score, G-Mean score (geometric mean of precision and recall), Precision score and Recall score. We do not apply accuracy as the positive and negative classes are not balanced.
Results: Experimental results are shown in Tab. 2. Notably, GPT-3.5 incorporated with our proposed prompting strategy even outperform the performance of GPT-4 incorporated with IO-prompting, Least-to-Most, CoT and CoT-SC, which have significantly lower recall scores, indicating their proneness to deception. Only Tree-of-Thoughts, which is benefited by its advantage in exploring various options, acquired the best results among all baselines, but is still defeated by our proposed method. Moreover, as we can see, for GPT-4 the performance of CoT-SC is even worse than CoT, which is supposed to be a specific case of CoT-SC without exploration. These results suggests that, when facing deceptive contents generated on purpose, existing incremental works’ improvement may not be so robust.
6 Discussions and Limitations
In summary, the proposed method has following advantages:
Comparison with IO-Prompting: Superiority in Expressive Power As we proved in Sec. 4, Compared to IO-prompting, our proposed method has stronger expressive power and thus can solve harder problems.
Comparison with CoT and EoT: Disentangling the task decomposition and task resolution Compared to the prompting family of CoT and EoT, our proposed method explicitly separate the task decomposition stage and task resolution stage. Therefore, we can acquire explicit decomposed sub-task rather than intermediate thoughts proposed during decoding. Consequently, we can explicitly enumerate all sub-tasks output by the decomposition module and avoid the model from missing important sub-tasks.
Comparison with LtM and Decomposed Prompting: Parallel Sub-task Handler and Sequential Sub-task Handler Similar as our proposed method, some program-guided prompting like LtM and Decomposed Prompting also explicitly separate the task decomposition stage and task resolution stage. However, they are mainly designed for multi-step reasoning for complex tasks. Thus, they sequentially tackle the sub-tasks and assembly the resolutions. As a result, they tend to follow the flow of the deceptive contents, leading to proneness to deceptive content. More details about this comparison can be found in Appendix A.4.
Although our proposed method DaC surpasses the baselines on the proposed tasks, it still has some limitations. The first issue is that the appliance scope of DaC is still limited. More specifically, CoT, EoT, LtM and DaC are based on different algorithmic paradigms, learning to different Appliance Scopes. As pointed out by Feng et al., CoT and LtM can be considered as a neural dynamic programming algorithm. Thus, CoT is more suitable for tasks that can be bridged to dynamic programming, such as multi-step question answering. Differently, EoT is based on exploration and search, which is more suitable for planning and search, such as Game of 24 (Yao et al., 2023). Our proposed method is based on Divide-and-Conquer algorithm. Thus, it is more suitable for tasks that can be decomposed to a series sub-tasks that are disjoint or only slightly overlapped. Our future work will focus on further expand the appliance scope of DaC to more areas like question answering.
7 Conclusions
In this paper, we proposed a novel Divide-and-Conquer program-guided problem solving strategy. To guide large language models to tackle tasks requiring large quantity of repetitive sub-tasks and/or deceptive contents, DaC disentangle the processes of task decomposition, sub-task resolution and resolution assembly. In this way, we avoid the intertwined processes from interrupting each other and alleviate the intermediate errors in generating the resolution path. The experimental results show that such improvement leads to better performance on a wide range of tasks such as large integer multiplication, hallucination detection and misinformation detection. Moreover, theoretic analysis reveals that our proposed method help guide LLM to expand the expressive power to surpass original Transformers.
Appendix A Appendix
A.1 Proof to Theorem 4.2
Before providing the proof, we first formally define how to organize the inputs (i.e., two 2-color trees) as a sequence. We assume that we acquire two trees $t_{p}$ of size $n$ and $t_{b}$ of size $m$. They are organized as two sequences of nodes with a random order. Each node has three variables: color, left child index, and right child index. If any child is null, then the index is filled with 0. Then we can organize them as as two sequences ${\bf X}_{p}\in\mathbb{R}^{n\times 3}$ and ${\bf X}_{b}\in\mathbb{R}^{n^{\prime}\times 3}$, where each item in the sequence is a vector of 3 dimensions. The first dimension is the index of the left child, the second dimension is the index of the right child, the third dimension is the color indicator (0 or 1). In addition, we have a root vector ${\bf r}$ with three dimensions. The first dimension is the index of the root node of $t_{p}$ (i.e., pointing to the root node of $t_{p}$) and the second is the index of the root node of $t_{b}$ (i.e., pointing to the root node of $t_{b}$). The third dimension of ${\bf r}$ is filled with 00 to make it have same dimension as the items in ${\bf X}_{p}$ and ${\bf X}_{b}$. This expression of trees is also called as pointer list encoding according to (Jenner et al., 2003). Note that in the following proof, we assume that all indices start from 1. Thus 0 is regarded as a NULL pointer.
Following the proof flow we provided in Sec. 4.2, we first provide the following divide-and-conquer algorithm that can solve the above problem:
Algorithm 3 Recursion Divide-and-Conquer Algorithm for 2-BSI $BSI({\bf r},{\bf X}_{p},{\bf X}_{b},m,t,d,f,w)$
0:  Inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, problem size metric function $f(\cdot)$, hyper-parameter $w$, merge function $m$, sub-task tackling function $t$, task decomposition function $d$
0:  A 0-1 indicator vector ${\bf v}$: if there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0.
1:  ${\bf r}_{l},{\bf r}_{r}\leftarrow d({\bf r},{\bf X}_{p},{\bf X}_{b})$
2:  for $i\in\{l,r\}$ do
3:     if $f({\bf r}_{i},{\bf X}_{p},{\bf X}_{b})>w$ then
4:        ${\bf v}_{i}\leftarrow BSI({\bf r}_{i},{\bf X}_{p},{\bf X}_{b},m,t,d,f,w)$
5:     else
6:        ${\bf v}_{i}\leftarrow t({\bf r}_{i},{\bf X}_{p},{\bf X}_{b})$
7:     end if
8:  end for
9:  Return $m({\bf r},{\bf X}_{p},{\bf X}_{b},{\bf v}_{l},{\bf v}_{r})$
Algorithm 4 Implementation of $d({\bf r},{\bf X}_{p},{\bf X}_{b})$ when the depth of the tree indicated by ${\bf r}$ is not longer than 2
0:  Inputs ${\bf r}\in\mathbb{R}^{3},{\bf X}_{p}\in\mathbb{R}^{n\times 3},{\bf X}_{b}\in\mathbb{R}^{n^{\prime}\times 3}$
0:  A 0-1 indicator vector ${\bf v}$: if there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0.
1:  ${\bf r}_{l}\leftarrow<{\bf X}_{p}[{\bf r}[1],2],{\bf r}[2],{\bf r}[3]>$
2:  ${\bf r}_{r}\leftarrow<{\bf X}_{p}[{\bf r}[1],3],{\bf r}[2],{\bf r}[3]>$
3:  Return ${\bf r}_{l},{\bf r}_{r}$
Algorithm 5 Implementation of $t({\bf r},{\bf X}_{p},{\bf X}_{b})$ when the depth of the tree indicated by ${\bf r}$ is not longer than 2
0:  Inputs ${\bf r}\in\mathbb{R}^{3},{\bf X}_{p}\in\mathbb{R}^{n\times 3},{\bf X}_{b}\in\mathbb{R}^{n^{\prime}\times 3}$
0:  A 0-1 indicator vector ${\bf v}$: if there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0.
1:  Initialize ${\bf v}$ as all q vector with a length of $n^{\prime}$
2:  if ${\bf r}[1]==0$ then
3:     Return ${\bf v}$
4:  end if
5:  for $i\in\{1,2,...,m\}$ do
6:     if ${\bf X}_{b}[i,3]!={\bf X}_{p}[{\bf r}[1],3]$ then
7:        ${\bf v}[i]\leftarrow 0$
8:     end if
9:  end for
10:  Return ${\bf v}$
Algorithm 6 Implementation of $m({\bf r},{\bf X}_{p},{\bf X}_{b},{\bf v}_{l},{\bf v}_{r})$
0:  Inputs ${\bf r}\in\mathbb{R}^{3},{\bf X}_{p}\in\mathbb{R}^{n\times 3},{\bf X}_{b}\in\mathbb{R}^{n^{\prime}\times 3},{\bf v}_{l}\in\mathbb{R}^{n},{\bf v}_{r}\in\mathbb{R}^{n}$
0:  A 0-1 indicator vector ${\bf v}$: if there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0.
1:  Initialize ${\bf v}$ as all 0 vector with a length of $n^{\prime}$
2:  if ${\bf r}[1]==0$ then
3:     Return ${\bf v}$
4:  end if
5:  for $i\in\{1,2,...,m\}$ do
6:     if ${\bf X}_{b}[i,3]=={\bf X}_{p}[{\bf r}[1],3]$ then
7:        if ${\bf v}_{l}[{\bf X}_{b}[i,1]]]==1$ and ${\bf v}_{r}[{\bf X}_{b}[i,2]]]==1$ then
8:           ${\bf v}[i]\leftarrow 1$
9:        else if ${\bf v}_{l}[{\bf X}_{b}[i,2]]]==1$ and ${\bf v}_{r}[{\bf X}_{b}[i,1]]]==1$ then
10:           ${\bf v}[i]\leftarrow 1$
11:        end if
12:     end if
13:  end for
14:  Return ${\bf v}$
The algorithm described above is a typical divide-and-conquer algorithm for solving rooted tree isomorphism. Its justification can be found in many textbooks introducing algorithms, such as Introduction to Algorithms (Cormen et al., 2022). Here we provide the detailed definition and implementation of problem size metric $f(\cdot)$, hyper-parameter $w$, merge function $m()$, sub-task tackling function $t(\cdot)$, task decomposition function $d(\cdot)$:
  • $w=1$, and $f({\bf r},{\bf X}_{p},{\bf X}_{b})$ is defined as the depth of the pattern tree $t_{p}$ indicated with root vector ${\bf r}$. Although precisely calculating $f({\bf r},{\bf X}_{p},{\bf X}_{b})$ is of $O(n)$, judging whether $f({\bf r},{\bf X}_{p},{\bf X}_{b})>1$ only require us to check whether the root node has child. If not, then return False.
  • $d({\bf r},{\bf X}_{p},{\bf X}_{b})={\bf r}_{l},{\bf r}_{r}$ returns two new root vectors ${\bf r}_{l},{\bf r}_{r}$. Both $r_{l},r_{r}$ have the same second and third dimension as ${\bf r}$. The ${\bf r}_{l}$’s first dimension is updated to be the index of the left child of the root node that ${\bf r}$ points to. The ${\bf r}_{r}$’s first dimension is updated to be the index of the right child of the root node that ${\bf r}$ points to. The updating function can be written as:
  • $t({\bf r},{\bf X}_{p},{\bf X}_{b})={\bf v}$ returns a 0-1 indicator vector ${\bf v}\in\mathbb{R}^{m}$ with the same length of the base tree size. If there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0. When the pattern tree’s depth is not higher than 1 (i.e., 1-node tree), $t({\bf r},{\bf X}_{p},{\bf X}_{b})$ is equivalent to output a 0-1 vector indicating the nodes in the base tree that have the same color of the root node of pattern tree. The implementation is provided in Alg. 5.
  • $m({\bf r},{\bf X}_{p},{\bf X}_{b},{\bf v}_{l},{\bf v}_{l})={\bf v}$ merge the results ${\bf v}_{l},{\bf v}_{l}$ to acquire a 0-1 indicator vector ${\bf v}\in\mathbb{R}^{m}$ with the same length of the base tree size. If there exists a subtree with node $i$ as root that is isomorphic with pattern tree $t_{p}$ defined with inputs ${\bf r},{\bf X}_{p},{\bf X}_{b}$, then the ${\bf v}[i]$ is 1. Otherwise, ${\bf v}[i]$ is 0. This function can be implemented by checking whether the pattern root’s children have a perfect match with each node’s children. Since each node has at most two children, checking the perfect match can be done in constant time. The implementation is provided in Alg. 6.
After providing the detailed implementation of the functions $d(\cdot),t(\cdot),m(\cdot)$, we are going to prove that there exists one unified transformer that can handle all these tasks with different prompts $d,t,m$. First, we will provide the following Lemma:
Lemma A.1.
Any fixed-size logic circuit that only contains multi-fan-in AND gates, multi-fan-in OR gates, NOT gates and has no recurrent structure can be precisely simulated by a multi-layer perceptron (MLP) with ReLU activation function and a width of $O(|Input|+|Circuit|)$ and a depth of $O(|Circuit|)$, where $|Input|$ denotes the size of input and $|Circuit|$ denotes the number of gates in the circuit.
Proof.
Assume that we are given a series of input pins with logic variable of 0 or 1, organized as a 0-1 vector ${\bf x}\in\mathbb{R}^{h}$. We first prove that all gates can be simulated by a two-layer perceptron. Then we can serialize all gates in the circuits and stack their corresponding 2-layer simulators accordingly to acquire a MLP simulator. An AND gate that take ${\bf x}$ as input can be simulated as:
$\text{AND}({\bf x})=\sigma(w_{A}{\bf x}-h+1)$ (1)
where $\sigma$ is the ReLU activation function, and $w_{A}$ is a weight vector with all dimensions equal to 1. If some dimensions of ${\bf x}$ are not the input of the gate, we can set the corresponding dimensions in the weight vector as 0 and adjust the $h$ as the input pin number. Similarly, an OR gate that take ${\bf x}$ as input can be simulated as:
$\text{OR}({\bf x})=1-\sigma(w_{O}{\bf x}+h+1)$ (2)
where $\sigma$ is the ReLU activation function, and $w_{O}$ is a weight vector with all dimensions equal to -1. A NOT gate is different, since it only takes one input pin. In such a case, we denote the index of the input pin as $i$, then we can simulate a NOT gate as:
$\text{NOT}({\bf x})=\sigma(w_{N}{\bf x}+1)$ (3)
where $w_{N}$ is a weight is a weight vector whose $i$-th dimension equals to -1 and all other dimensions equal to 0. Also, since the $x$ is a 0-1 vector, the activation function is equivalent to a identical function to $x$:
${\bf x}=\sigma({\bf x})$ (4)
To construct a MLP that can simulate a fixed-size logic circuit without recurrent structure, we apply the circuit serialization in (Merrill & Sabharwal, 2023b) which order the gates based on topological order. In this way, we can represent the circuit as a sequence GATE[1], GATE[2], GATE[3],…,GATE[L], where each GATE[i]’s input only contains the output of the previous gates and the original input ${\bf x}$. Therefore, we can construct a $2L$-layer MLP base on the above serialization. Specifically, the $2i$-th and $12i+$ 1-th layers of the MLP will simulate the $GATE[i]$ as well as copy all previous inputs with activation function and concatenate them together. This can be done by concatenate an identical matrix on the GATE’s weight vector ($w_{A}$, $w_{O}$ or $w_{N}$). In this way, we can construct a MLP that precisely simulate the circuit. Since every time we concatenate the output of a gate with the input of it, the input dimension number of the final layer can be bounded by $O(|{\bf x}|+L)$. In the worst case, for a circuit of size $L$, we needs $2L$ layers to precisely simulate it. However, in many cases, a lot of gates in the circuits can be run parallelly. In such cases, the MLP could be much more shallow.
Now, we can start to prove our main theorem:
Theorem A.2.
There exists a log-precision transformer with fixed depth and hidden dimension that can solve the 2-BSI of any size with fixed-length prompt $m$ (for merge), $t$ (for sub-task tackling) and $d$ (for task decomposition).
Proof.
We prove this theorem by constructing a Transformer that can tackle this problem. First we define how to organize the input given ${\bf r},{\bf X_{p}},{\bf X_{b}}$ and the prompt. Specifically, we construct a feature sequence ${\bf X}\in\mathbb{R}^{(3+n+n^{\prime})\times 7}$. Each item in this sequence is a feature of 7 dimensions, indicating a token. The first two dimensions indicate whether the token is a prompt (’00’), a root vector (’01’), a pattern tree node (’10’), or a base tree node (’11’). The third to fifth dimensions carries the information about the token. For a prompt token, ’100’ indicates merge prompt $m$, ’010’ indicates sub-task tackling prompt $t$, and ’001’ indicates task decomposition prompt $d$. For other cases, these three dimensions are with the same formula as the three dimensions in ${\bf r},{\bf X}_{p},{\bf X}_{b}$. The rest two dimensions are allocated specifically for the merge function $m(\cdot)$ to store ${\bf v}_{l}$ and ${\bf v}_{r}$. More specifically, for the feature of token indicating the $i$-th base tree node, its sixth dimension is ${\bf v}_{l}[i]$ and its seventh dimension is ${\bf v}_{r}[i]$. For other tokens, these two dimensions are filled with 0. In ${\bf X}[1]$ we store the prompt token. In ${\bf X}[2]$ and ${\bf X}[3]$ we store the input root vector ${\bf r}$ duplicately. We store the same token twice so that we can tackle ${\bf r}_{l}$ and ${\bf r}_{r}$ separately. To separate this two token, we use the last dimension, which was padded as 0 in ${\bf r}$, to distinguish them. ${\bf X}[2,5]$ is set as 0 and ${\bf X}[3,5]$ is set as 1. From ${\bf X}[4]$ to ${\bf X}[3+n]$, we store ${\bf X}_{p}$. From ${\bf X}[4+n]$ to ${\bf X}[3+n+n^{\prime}]$, we store ${\bf X}_{b}$. For all node indices of pattern tree, we add them by 3. For all node indices of base tree, we add them by 3+n, so that the indices can be applied to directly retrieve the positional embeddings.
Algorithm 7 Logic circuit for MLP of the second Transformer layer
0:  Input feature ${\bf x}^{\prime\prime}\in\mathbb{R}^{42}$
0:  Output feature ${\bf y}\in\mathbb{R}^{7}$
1:  ${\bf y}\leftarrow{\bf x}^{\prime\prime}[1:7]$Initialize ${\bf y}$}
2:  if ${\bf x}^{\prime\prime}[1:2]==00$ or ${\bf x}^{\prime\prime}[1:2]==10${Prompt Token or Pattern Tree Node} then
3:     Return ${\bf y}$
4:  else if ${\bf x}^{\prime\prime}[1:2]==01$Root Vector Token} then
5:     if ${\bf x}^{\prime\prime}[24:26]==001${Prompt is $d$then
6:        if ${\bf x}^{\prime\prime}[5]==0$ then
7:           ${\bf y}[3]\leftarrow{\bf x}^{\prime\prime}[10]$get ${\bf r}_{l}$, similar as line 1 in Alg. 4}
8:        else if ${\bf x}^{\prime\prime}[5]==1$ then
9:           ${\bf y}[3]\leftarrow{\bf x}^{\prime\prime}[11]$get ${\bf r}_{r}$, similar as line 2 in Alg. 4}
10:        end if
11:     end if
12:  else if ${\bf x}^{\prime\prime}[1:2]==11$Base Tree Node Token} then
13:     if ${\bf x}^{\prime\prime}[24:26]==010${Prompt is $t$then
14:        if ${\bf x}^{\prime\prime}[40]=={\bf x}^{\prime\prime}[5]${Line 6 in Alg. 5then
15:           ${\bf y}[5]\leftarrow 1$
16:        else
17:           ${\bf y}[5]\leftarrow 0$
18:        end if
19:     else if ${\bf x}^{\prime\prime}[24:26]==100${Prompt is $m$then
20:        if ${\bf x}^{\prime\prime}[13]==1$ and ${\bf x}^{\prime\prime}[21]==1$Line 7 in Alg. 6then
21:           ${\bf y}[5]\leftarrow 1$
22:        else if ${\bf x}^{\prime\prime}[14]==1$ and ${\bf x}^{\prime\prime}[20]==1${Line 9 in Alg. 6then
23:           ${\bf y}[5]\leftarrow 1$
24:        else
25:           ${\bf y}[5]\leftarrow 0$
26:        end if
27:     end if
28:  end if
After preparing the inputs, we start to construct our Transformer. The transformer first attach the position index for each token (positional embedding). After that, the inputs are forwarded into a transformer with depth of 2. Each transformer layer contains a multi-head attention layer followed by a MLP. As proved by (Merrill & Sabharwal, 2023b; Feng et al., 2023), the attention layer of Transformer can retrieve the feature of tokens whose positional embeddings satisfy specific conditions. For multi-head attention, different heads can retrieve tokens with different conditions. In the following construction, we will use this conclusion to construct attention heads with different functions.
In the first Transformer layer, the function of each attention head is defined as:
  • Head 1 only attends to the token itself to store $X[i]$ for token $i$.
  • Head 2 attends to the token with a positional embedding matches the ${\bf X}[i,3]$ and copy this token’s 5-dimension feature. For tree node tokens, this head’s job is to retrieve the feature of ${\bf X}[i]$’s left child. For root vector tokens, this head’s job is to retrieve the feature of pattern tree root node. For the first token (prompt token), this head’s retrieved feature will not be applied in the afterwards layers and thus does not influence the correctness of the model.
  • Similar as Head 2, Head 3 attends to the token with a positional embedding matches the ${\bf X}[i,4]$ and copy this token’s 5-dimension feature. This head’s job is to retrieve the feature of ${\bf X}[i]$’s right child. For root vector tokens, this head’s job is to retrieve the feature of base tree root node.
  • Head 4 attends to the first token (prompt token) and copy this token’s 7-dimension feature. This head’s job is to retrieve the prompt indicator.
  • Head 5 attends to the second token (root token) and copy this token’s 7-dimension feature. This head’s job is to retrieve the root information.
With the above 5 heads, the attention layer will output a 35-dimension feature for each token. We denote these features as ${\bf X}^{\prime}\in\mathbb{R}^{(3+n+n^{\prime})\times 35}$. After that, these features are forwarded into a MLP fitting identical mapping to acquire the input features for the second Transformer layer.
In the second Transformer layer, the function of each attention head is defined as:
  • Head 1 only attends to the token itself to store $X^{\prime}[i]$ for token $i$.
  • Head 2 attends to the token with a positional embedding matches the ${\bf X}^{\prime}[i,31]$ and copy this token’s 1-7 dimension features (${\bf X}^{\prime}[{\bf X}^{\prime}[i,31],1:7]$). This head’s job is to broadcast the feature of the pattern tree root node to every token.
With the above 2 heads, the attention layer will output a 42-dimension feature for each token. We denote these features as ${\bf X}^{\prime\prime}\in\mathbb{R}^{(3+n+n^{\prime})\times 42}$. For root vector token, only the features from head 1 and head 4 are useful. For base tree node tokens, all 42 dimensions are useful. Then each token’s feature are parallely forwarded into a MLP. We will use this MLP to fit the logical circuit described in Alg. 7. The function of Alg. 7 is to aggregate the functions of $m(\cdot),t(\cdot),d(\cdot)$ together and assign the correct value based on the prompt indicator. In Alg. 7, all operations are AND, OR, NOT, SELECTOR, and ASSIGN and there is not loop. Thus, it is a static logical circuit and can be implemented with multi-fan-in AND, OR, NOT gates. Thus, it can be precisely simulated by a MLP according to our Lemma A.1.
After acquiring the ${\bf y}\in\mathbb{R}^{7}$ for each token, we can organize them as a feature sequence ${\bf Y}\in\mathbb{R}^{(3+n+n^{\prime})\times 7}$. When the prompt is $d$, we return ${\bf Y}[2,3:5]$ as ${\bf r}_{l}$ and ${\bf Y}[3,3:5]$ as ${\bf r}_{r}$. If the prompt is $t$ or $m$, then we can output ${\bf Y}[3+n+1:3+n+n^{\prime},5]$ as the expected ${\bf v}$.
A.2 Prompting Details of DaC
Multiplication of Long Integers: Suppose we have two $2n$-digit numbers $AB$ and $CD$, where $A,B,C,D$ are all $n$-digit numbers. Then we can break $AB\times CD$ as $(A\times C\times 10^{2n})+(A\times D\times 10^{n})+(B\times C\times 10^{n})+(B\times D)$, where the calculation in each bracket pair is disjoint with others bracket pairs. We only need to compute the results of multiplication in each bracket pair parallelly and then merge all of them with addition:
Decomposer Prompt $d$: Please split the string a from the middle as two separated strings. The lengths of the two separated strings should be as close as possible. Please only return the two strings separated by a comma and do not return anything else.
Sub-task Tackling Prompt $t$: (1)Please compute $a*b$. (2) Please only return the final results and do not return anything else (ensure disentangled-sub-process principle).
Merge Prompt $m$: Please compute $x=a*10^{2n}+b*10^{n}$ and $y=c*10^{n}+d$. Based on the above calculation, please compute $x+y$ carefully step by step.
Hallucination Detection in Long Context: We divide the summary to sentences. After that, we paralelly verify the sentences. Finally, we merge the verification to each sentence:
Decomposer Prompt $d$: Please help me segment the following paragraph as sentences. The separated sentence should be output as: #Statement 1#: …#Statement 2#: …Do not say anything else. Just return the statements in the given format.
Sub-task Tackling Prompt $t$: I want you to act as a factual contradiction checker. You are given a set of statements and a document. Among the statements, there might be one or more statement that contains contradictions with the document. Please find the problematic statement if it exist by analyzing the statements one by one. For each statement, please make a choice:
  • A: The statement is totally aligned with the document for sure.
  • B: The statement contradicts with the document.
Merge Prompt $m$: Based on the above analysis, please tell me, does any statement above contain contradiction with the document?.
Fact-Verification for Misinformation Detection: Similar as hallucination detection, we divide the summary to sentences. After that, we paralelly verify the sentences. Finally, we merge the verification to each sentence. Thus, our decomposer prompt and sub-task tackling prompt are the same as hallucination detection. The only difference is the merge prompt.
Merge Prompt $m$: If we connect the above statements to be a news article, based on the above analyzation, please answer me: Is there any contradiction between the document and the article?
Figure 5: Comparison of Least-to-Most (LtM) Prompting and Decomposed Prompting (DeP).
A.3 Decomposed Prompting and Least to Most
Least-to-Most (LtM) Prompting (Zhou et al., 2022) and Decomposed Prompting (Khot et al., 2022) are two similar works to our work. They both propose to explicitly prompt the LLM to decompose the task as a series of sub-tasks and sequentially tackle them. In Fig .2, we merge these two methods. Here, we will provide more detailed comparison of them, which is shown in Fig. 5. Decomposed Prompting can regarded as a upgraded version of LtM. It introduces special notations into the prompt to represent program states so that when sequentially tackling the sub-tasks, it can call heterogeneous modules to tackle them. Such design enable the LLM to call external programs (e.g., retrieval documents on WikiPedia and program based calculator) and/or itself (i.e., recursion). Such design endows it stronger expressive power and increases the compositional generalization ability of LLMs in different areas, such as symbolic manipulation and multi-hop QA (Khot et al., 2022). Also, it endows LLM the ability to do open-domain QA by retrieving from external knowledge base.
A.4 More Discussions on Sequential Sub-task Tackling and Parallel Sub-task Tackling
Figure 6: Toy example of Sequential Sub-task Tackling and Parallel Sub-task Tackling in long integer multiplication
Figure 7: Toy example of Sequential Sub-task Tackling and Parallel Sub-task Tackling in hallucination detection
Sequential Sub-task Tackling and Parallel Sub-task Tackling are two different paradigm in decomposing complex tasks as sub-task to tackle. The first one decompose a complex tasks as a series of sub-tasks. In this series, each sub-task relies on the previous one’s output as input or context. The second one decompose a complex tasks as a set of sub-tasks, each of which does not rely on others. Two examples for multiplication and hallucination detection are provided in Fig 6 and 7